OpenStudy (loser66):

What is the real and imaginary parts of \(i^n, 2\leq n\leq 8\) Please, help

OpenStudy (loser66):

What I don't get is when n is odd, like 3, 7, then \(i^3 = -i\) and \(i^7 = -i\) also. But for n =5, we have \(i^5 = i\) , hence, we can classify n is odd as a common case to determine Real and Imaginary, right?

OpenStudy (loser66):

Similarly to n is even, if n =2, then \(i^2 =-1\), but if n =4, 8 \(i^4 =1 ~~and~~ i^8 =1 \) We have to break it into many cases and I don't want it. Is there any way to combine them to general cases? Because if the limits are \(2\leq n\leq 100\), how can I list all of the cases out?

OpenStudy (anonymous):

what's i?\[i=\sqrt{-1}\] which is totally imaginary.... \[i ^{2}=(\sqrt{-1})^{2} =-1\] but \[i ^{4} = (\sqrt{-1})^{4}= i^{2}*i ^{2}= (-1)*(-1)= 1\]

OpenStudy (anonymous):

if u get this, u can understand the whole concept.........

OpenStudy (loser66):

@lall Thanks for the answer. I got it. All I want is to classify the cases to not list them up.

OpenStudy (anonymous):

\[i ^{5}=i ^{4+1}= i ^{4}*i ^{1}= 1* i=i\]

OpenStudy (loser66):

How about this question: Find Real and Imaginary part of \(\left (\dfrac{1+i}{\sqrt2}\right )^n, 2\leq n\leq 8\)

OpenStudy (loser66):

Again, I don't want to list the cases. Please, help me or tell me "No choice for you" :)

OpenStudy (irishboy123):

tried the exponential form? so \(\large e^{i \ n\frac{\pi}{4}}\) is handy notation

OpenStudy (loser66):

Is it not that it is NOT a cis problem? If we convert it to cis, then we have to convert back to complex to find R and Im, right?

ganeshie8 (ganeshie8):

\[i^n\equiv i^{n\pmod{4}}\]

OpenStudy (phi):

I assume you noticed for the first question \[ i^n= i^{n\ mod\ 4} \] for the second, it is easier to convert to polar (or cis ) form, do the calculation, then convert back to rectangular

OpenStudy (phi):

also, notice the answers will be equally spaced on the unit circle, starting at i

ganeshie8 (ganeshie8):

i feel the second problem can be done easily by looking at geometry : notice that the given complex number lies exactly halfway between real and imaginary axes and recall the fact that angles add up when you multiply two complex numbers

OpenStudy (loser66):

Show me, please.

ganeshie8 (ganeshie8):

let \(z=\dfrac{1+i}{\sqrt2}\) it is of unit length, so each power increments the angle by 45 degrees but the length stays same : 

ganeshie8 (ganeshie8):

all the red lines are of unit length ^

OpenStudy (loser66):

oh, you convert to cis, right?

OpenStudy (phi):

Irish posted the form you should use

OpenStudy (loser66):

I got it. Thanks a lot, friends :)

ganeshie8 (ganeshie8):

np, you must have noticed \(\left (\dfrac{1+i}{\sqrt2}\right )^n, 1\leq n\leq 8\) give you the eight roots of the equation \(z^8-1=0\) and they also form a group under regular multiplication of complex numbers

OpenStudy (loser66):

Yes, Sir.

OpenStudy (loser66):

Question @ganeshie8 I have to find |z|, I confused. \(z= (1+i)^6\) My logic: \(|1+i|= \sqrt 2\rightarrow |1+i|^6 =8\) If we convert to cis, like what you did, then  What is wrong?